Sale of Equipment Equipment was acquired at the beginning of the year at a cost of $33,250. The equipment was depreciated using the double-declining-balance method based orn an estimated useful life of ten years and an estimated residual value of $650. a. What was the depreciation for the first year? b. Assuming the equipment was sold at the end of year 2 for $7,680, determine the gain or loss on the sale of the equipment. c. Journalize the entry to record the sale. If an amount box does not require an entry, leave it blank.
Answer #1
a
Depreciation for first year = 33250*20%= $6650
b
Depreciation for second year =(33250-6650)*20%= $5320
Book value at the date of sale = 33250-(6650+5320)= $21280
Loss on sale of Equipment = 21280-7680= $13600
c
Cash 7680
Accumulated depreciation-Equipment 11970
Loss on sale of Equipment 13600
       Equipment 33250